ADNAN MUHAMMED

Grade 12,

A+B=π/2 and B+C=A, then tan A=? Please reply as soon as possible.

A+B=π/2 and B+C=A, then tan A=?
Please reply as soon as possible. 

Grade:

1 Answers

Rohit verma
5 years ago
  1. TanA=tan pi-B
  2.  TanA= cotB
  3.  C=A-B
  4. TanC = Tan (A-B)
  5. Tanc=tanA-tanB/1+tanAtanB
  6. TanA=cotB
  7.  Hence tanC = tan A - tanB/1+1
  8. 2TanC=tanA-TanB
  9. Hence tTanA=2TanC+tanB
  10. 2tanC=tanA -tan